Overheads are budgeted to be £100,000 for the year. Labour hours are 50,000, Machine hours 20,000 and units produced 25,000. What would the blanket overhead rate be using the following three bases of allocation:

(a) Labour hours

(b) Machine hours

(c) Units
Answer #1

When a single overhead rate is computed for the factory as a whole it is known as single or blanket or plant wide rate.

Blanket Rate = Overhead cost for the entire factory/Total quantum of the base selected

Blanket rate is applied in small concerns where only one product is manufactured on a continuous basis or where all the products pass through all the operations or departments and the incidence of overhead is uniform on all the departments. These rates are easy to compute and require less clerical cost but have a very limited use.

Total Overheads = 100,000

a)Blanket Overhead Rate = Total Overheads/Total Labour hours

Blanket Overhead Rate = 100000/50000

                                          = £ 2 per labour hour

b)

Blanket Overhead Rate = Total Overheads/Total Machine hours

                                           = 100000/20000

                                           = £ 5 per machine hour

c)

Blanket Overhead Rate = Total Overheads/Total units produced

                                            = 100000/25000

                                           = £ 4 per unit